OpenStudy (rational):

Let N = 123123123... (300 digits). Find the remainder when N is divided by 504

OpenStudy (ikram002p):

\(\sum_{n=0}^{99}123\times 10^{3n} \mod 504\)

OpenStudy (ikram002p):

And 504= 3*2^3*23

OpenStudy (rational):

504 = 7*8*9

OpenStudy (ikram002p):

Ok so we need find construction of 10^3n mod 504

OpenStudy (ikram002p):

Lol that was stupid xD

OpenStudy (ikram002p):

Ok there is a pattern 123+ 123(-8+64-8+64+...)

OpenStudy (ikram002p):

123+123(49*56)+123*-8. Mod 504

OpenStudy (ikram002p):

So we would have 123(1+2744-8) mod 504 123(1+224-8) mod 504 123*217 mod 504

OpenStudy (ikram002p):

I would say remainder is 483 Hope it's right calculating mentality frustrated me xD

OpenStudy (ikram002p):

I would like also to see fancy theorems or algorithms it would be fun

OpenStudy (rational):

483 is correct! im actually looking for a more elementary method without using so much of number theory as this particular question was asked in one of the highschool level competitions..

OpenStudy (empty):

I was looking for a way to combine the fact that \[504=7*8*9=n*(n+1)(n+2)=1n^3+3n^2+2n \text{ for n=7}\] into something that exploits that 1,3,2 to somehow do magical things to the 1,2,3 while avoiding touching the tempting geometric series.

OpenStudy (empty):

Or maybe it's just a coincidence... haha... I sort of took that 504=7*8*9 to be almost a type of hint.

OpenStudy (rational):

Ahh that might give some easy criterian for divisibility test for 7, not so sure, still thinking..

Parth (parthkohli):

I think I got it.

OpenStudy (rational):

thats a neat approach! let me see if i understand it correctly.. for n=5, im getting remainder =291 ? http://www.wolframalpha.com/input/?i=123*10%5E%283*%285-1%29%29+%2B+123123123123+mod+504

Parth (parthkohli):

Oooh, oops. I hate my calculator.

OpenStudy (rational):

this is easy with chinese remainder theorem i just don't understand why they give these problems in highschool level competitions

Parth (parthkohli):

Sorry, I got everything wrong. That can never be zero.

Parth (parthkohli):

Which competition?

OpenStudy (rational):

this question is everywhere. i saw this for the first time on brainly last night, google gave me million+ results.. mostly from various competitions in india

OpenStudy (rational):

but i couldn't find any satisficatory solution anywhere online

Parth (parthkohli):

lol, brainly? seriously? D:

OpenStudy (rational):

Haha thats the reason i need to dumb it down to highschool level, otherwise they are deleting my answers

Parth (parthkohli):

That's foolish. If a method provides a legitimate solution, no matter how incomprehensible it is, it should be left there. Sounds like a dumb reason to delete answers.

OpenStudy (ikram002p):

Watching the hundred feet journey, I have some other methods though

OpenStudy (rational):

`Reason Hi! I'm sorry but your answer lacks some vital details and therefore, it is being deleted. Please elaborate your answer further and repost it.`

OpenStudy (rational):

please post after the movie ikram

OpenStudy (ikram002p):

i thought it was that easy thats why i used less NT hmmm other method is deleling with it as series

OpenStudy (ikram002p):

hmm no not interesting anymore.. i hate hight school anyway

OpenStudy (ikram002p):

dealing*

OpenStudy (lalipapa123):

hmm

OpenStudy (rational):

hey are you from brainly

OpenStudy (lalipapa123):

yeah..

OpenStudy (rational):

welcome! :)

OpenStudy (lalipapa123):

thnks

OpenStudy (ikram002p):

what is brainly ?

OpenStudy (ikram002p):

seems i have account there :O since when XD

OpenStudy (ikram002p):

http://brainly.com/profile/ikram002p-98608

OpenStudy (rational):

yes that site i think it forces you to create account to see answer to a problem

OpenStudy (ikram002p):

I see